OpenStudy (anonymous):

Help? Rewrite with only sin x and cos x. sin 2x - cos 2x

OpenStudy (anonymous):

@jim_thompson5910 @mathmale @Mertsj Someone please help me.. :(

OpenStudy (mertsj):

\[\sin ^2x-\cos ^2x \] or\[\sin (2x)-\cos (2x)\]

OpenStudy (anonymous):

That's what I thought, too. But these are my choices: 2 sinx cosx - 1 + 2 sin^2x 2 sin x cos2x - 1 + 2 sin^2x 2 sin x cos2x - sin x + 1 - 2 sin^2x 2 sin x cos2x - 1 - 2 sin^2x

OpenStudy (anonymous):

Come back.. I'm so lost.. :'c

OpenStudy (anonymous):

@RadEn @satellite73 @whpalmer4

OpenStudy (raden):

use these identities : sin(2x) = 2sin(x)cos(x) cos(2x) = 1 - 2sin^2 (x)

OpenStudy (whpalmer4):

\[\sin (2x)-\cos(2x)\]appears to be the correct interpretation.

OpenStudy (anonymous):

But none of my answer choices match up to that...

OpenStudy (mertsj):

Would it be too much trouble for you to tell which interpretation of the problem is the correct one?

OpenStudy (anonymous):

What do you mean?

OpenStudy (mertsj):

Would it be too much trouble for you to read the first post I wrote?

OpenStudy (anonymous):

Actually, I think the way Raden described it gives me answer choice A.

OpenStudy (whpalmer4):

Yes, if you interpret it as \(\sin(2x)-\cos(2x)\) then answer A is correct.

OpenStudy (anonymous):

Oh! That's what you were saying. Makes sooooo much more sense now. \[\sin 2x - \cos 2x\] It's not exponents..

OpenStudy (whpalmer4):

A lot of collective time is wasted working problems that people haven't clearly stated, and it's very helpful if you either a) state it unambiguously from the outset, or b) answer questions from people asking "do you mean this, or that" :-)

OpenStudy (whpalmer4):

the problem is that people copy and paste, or don't know any better, and things like \(\cos^2x\) in their problem become \(\cos2x\) on our screens...

OpenStudy (anonymous):

My bad. I didn't do it in the equation thing because the way I put it the first time was right. If it has exponents, I will use this thing--> x^2.

OpenStudy (whpalmer4):

I'm just explaining why my colleague might come across as being a bit irritated...you wouldn't like him when he's irritated, so avoid that :-)

OpenStudy (anonymous):

Haha. Sorry all. I thought I was clear. Thanks for all of your wonderful help. God bless you.

OpenStudy (whpalmer4):

In general, look at what you write, and ask yourself, "is there any possible way that someone could interpret this in a different fashion than I intend?" If so, add parentheses or use the equation editor (I think most of us assume that if you use the EE, you mean what you write, but no doubt I'll soon waste 15 minutes working a problem where that isn't true :-)

OpenStudy (whpalmer4):

The less time we waste on this, the more time we have to actually help :-)

OpenStudy (anonymous):

Haha. :)